If a bowling ball is 8 inches in diameter, what is it's volume to the nearest cubic inch ?
Formula : V= 4/3 (3.14)R3 ----- THAT IS "V" equals pie times r to the 3rd power
I have never worked with any problems like this and just really wanted to see someone work it out !
: If a bowling ball is 8 inches in diameter, what is it's volume to the nearest cubic inch ?
Formula : V= 4/3 (3.14)R3 ----- THAT IS "V" equals pie times r to the 3rd power
I have never worked with any problems like this and just really wanted to see someone work it out !

Answer by nerdybill(419)

The problem gives you the volume of a sphere:
V= 4/3 (3.14)R^3
where
V is the volume
R is the radius
.
Also, in the problem, they said that the bowling ball is 8 inches in diameter.
Since, half the diameter is the radius we have:
R = (1/2)(8)= 4 inches
.
Now we simply plug it into our equation:
V= 4/3 (3.14)(4)^3
First, multiplying the 4 and 3.14 we get:
V= (12.56/3)(4)^3
Dividing 12.56 by 3:
V= (4.187)(4)^3
Then we resolve 4^3:
V= (4.187)(64)
Finally, we get:
V= 267.947 cubic inches

Question 149760: If a bowling ball is 8 inches in diameter, what is it's volume to the nearest cubic inch ?
Formula : V= 4/3 (3.14)R3 ----- THAT IS "V" equals pie times r to the 3rd power
I have never worked with any problems like this and just really wanted to see someone work it out !
: If a bowling ball is 8 inches in diameter, what is it's volume to the nearest cubic inch ?
Formula : V= 4/3 (3.14)R3 ----- THAT IS "V" equals pie times r to the 3rd power
I have never worked with any problems like this and just really wanted to see someone work it out !

Answer by Earlsdon(3472)

The volume of a bowling ball with 8" diameter:
V = (4/3)(pi)r^3 r = half the diameter. = 8/2 = 4"
V = (4/3)(3.14)(4^3)
V = (12.56/3)(64)
V = 267.9466 Round to nearest cubic inch...
V = 268cubic inches.